OpenStudy (anonymous):

what is the value of the expression sqrt(27)+sqrt(32)-sqrt(75)

OpenStudy (anonymous):

\[\sqrt(27)+\sqrt(32)-\sqrt(75)\]

OpenStudy (anonymous):

@zaynahf if your not busy can you help me out?

OpenStudy (anonymous):

Sure, i can! Do you know how to simplify radicals?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

>< ---

OpenStudy (anonymous):

To simplify radicals, you have to try to find a perfect square that fits perfectly into the number. So for the first one: \[\sqrt{27}\] the biggest perfect square that can fit into it is 9. So, the inside would simplify to be: \[\sqrt{9 \times 3} = 3\sqrt{3}\]

OpenStudy (anonymous):

so i can get \[\sqrt16X2 = 4\sqrt2 ?\]

OpenStudy (anonymous):

Yeah! Now try the third one

OpenStudy (anonymous):

\[\sqrt75 = \sqrt5X15 = 5\sqrt5?\]

OpenStudy (anonymous):

No, since neither 15 nor 5 are perfect squares. Here are the perfect squares: 1,4,9,16,25,36,49,64,81,100....

OpenStudy (anonymous):

so \[aaaannnnnd`im lost\]

OpenStudy (anonymous):

lol i'll guide you through

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Ok since you want to get the biggest one that goes into there, start from the top of the list. Try seeing if 100 goes into 75 (obviously not), so just start from the smallest number after 75. 64,36,25,16,4,1. When you find the one that goes into 72, let me know.

OpenStudy (anonymous):

64

OpenStudy (anonymous):

64 goes into 75 evenly?

OpenStudy (anonymous):

oh no 25

OpenStudy (anonymous):

Good! Now what times 25 is 75?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

So thats what goes under the radical: \[\sqrt{75}= \sqrt{25 \times 3}= 5\sqrt{3}\] Right?

OpenStudy (anonymous):

ok what do i do now multiply them together?

OpenStudy (anonymous):

\[3\sqrt{3}+4\sqrt{2}-5\sqrt{3}\] Is that what you have so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

There are only 2 with the same base radical, so you can only simplify those

OpenStudy (anonymous):

so i get \[8\sqrt3+ 4\sqrt2\]

OpenStudy (anonymous):

Remember, it's 3-5, not 3+5

OpenStudy (anonymous):

-2√3+4√2

OpenStudy (anonymous):

is that my final answer?

OpenStudy (anonymous):

Perfect

OpenStudy (anonymous):

Yes it is

OpenStudy (anonymous):

thnxs so much ^_^ now i can sleep ... eh maybe

OpenStudy (anonymous):

No problem! Let me know if you need other help